A mail-order computer business has six telephone lines. Let X denote the number of lines in use at a specified time.

Suppose the pmf of X is as given in the following table. x 0 1 2 3 4 5 6

p(x) 0.10  0.15 0.20 0.25 0.20 0.06 0.04

(a) (4 points) Calculate P(2 < x ≤ 5).

(b) (4 points) Write down the cdf of X.
Answer #1

a)P(2<X≤5)=P(X=3)+P(X=4)+P(X=5)=0.25+0.2+0.06=0.51

b) cdf is given by F(x)=0 if x<0

0.1 if 0≤x<1

0.25 if 1≤x<2

0.45 if 2≤x<3

0.7 if 3≤x<4

0.9 if 4≤x<5

0.96 if 5≤x<6

1 if x≥6

If you do not get anything in this solution, please put a comment and I will help you out. Do not give a downvote instantly. It is a humble request. If you like my answer, please give an upvote.